OpenStudy (anonymous):

what algebraic property are these? 1.)5x + 24 = 59 FOR THIS ONE I PUT Associative Property 2.)5x + 0 = 35 FOR THIS ONE I PUT subtraction Property of Equality 3.)5x = 35 FOR THIS ONE I PUT Division Property of Equality 4.) x = 7 FOR THIS ONE I PUT reflexive property AM I DOING THIS CORRECTLY?

OpenStudy (anonymous):

I think the option 1 is wrong ( I think)

Directrix (directrix):

Is this --> 1.)5x + 24 = 59 the given equation? If that is the beginning point of the problem, then why did you pair that statment with Associative Property? I think your reasons for steps 3 and 4 are wrong. Did you post all the steps? @klawdeeahh

OpenStudy (anonymous):

Given: 5(x + 3) + 9 =59 Prove: x = 7 Statements 1) 5(x + 3) + 9 = 59 2) 5x + 15 + 9 = 59 3) 5x + 24 = 59 4) 5x + 0 = 35 5) 5x = 35 6) x = 7 A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

Directrix (directrix):

Statements 1) 5(x + 3) + 9 = 59 -----------> Given 2) 5x + 15 + 9 = 59 3) 5x + 24 = 59 4) 5x + 0 = 35 5) 5x = 35 6) x = 7

OpenStudy (anonymous):

From what I can see, looks like you have 1, 3, and 4 wrong from your original post..

OpenStudy (anonymous):

5(x+3) + 9 = (5*x)+(5*3) + 9 = 5x+15 + 9 you're distributing the 5 on the outside of the parentheses to all terms within through multiplication, so that one's distributive property.

OpenStudy (anonymous):

next one you're using addition to find a number equal to 15+9 -- so addition property of equality

OpenStudy (anonymous):

subtracting 24 from both sides to keep both sides equal -- subtraction property of equality

OpenStudy (anonymous):

#1 i know its Given #2 distribitive #3 not quiet sure #4 additive identity? #5 mutiplicative identity? #6 Symmetric Property?

OpenStudy (anonymous):

5x+0 = 5x additive identity, any number plus zero is itself

OpenStudy (anonymous):

1 Given 2 Distributive 3 Addition Property of Equality 4 Subtraction Property of Equality 5 Additive Identity 6 Division Property of Equality

OpenStudy (anonymous):

If you want explanations i'll try my best

Directrix (directrix):

Statements 1) 5(x + 3) + 9 = 59 -----------> Given 2) 5x + 15 + 9 = 59 -----------> Distributive Property 3) 5x + 24 = 59 -----------> Substitution Property (24 for 15 + 9) 4) 5x + 0 = 35 ------------> Subtraction Property 5) 5x = 35 ------------> Additive Identity 6) x = 7 ------------> Division Property of Equality @klawdeeahh This "proof" is, in my opinion, missing some steps as well as possible reasons. Regardless, what I have written here is my best attempt to help work a poorly created question. Therefore, I do not claim that this work is correct.